OpenStudy (anonymous):

determine the solutions of 3lx+6l=9x-6 (absolute value)

OpenStudy (anonymous):

|x+6|=3x-2 "left side of the abs.v." wont intersect 3x-2 so: x+6=3x-2 2x=8 x=4

OpenStudy (anonymous):